OpenStudy (anonymous):

Find the inverse of m(x)=√ x+1/x, x>0

OpenStudy (anonymous):

Is it one to one though?

OpenStudy (anonymous):

\[ \lim_{x\to 0}m(x) = \infty \\ \lim_{x\to \infty}m(x) = \infty \\ \]It's not one to one, so it doesn't have an inverse.

OpenStudy (anonymous):

books says the answer is m^-1(x)=1/x^2-1

OpenStudy (anonymous):

Then either you wrote the equation wrong or the book is wrong.

OpenStudy (anonymous):

\[ \large m^{-1}(m(x)) = \frac{1}{(\sqrt{x}-\frac{1}{x})^2}-1 \neq x \]